1. Suppose that the total utility function of a consumer is given by TU(x,y) = 6XY, and the prices of X and Y are 1 birr and 2 birr per unit, respectively. If the income of the consumer is birr 600 birr and if he/she spends all of his/her income on the consumption of commodities of X and Y, find the optimum amount of X and Y that the consumer will consume at equilibrium.
1
Expert's answer
2020-01-24T09:46:51-0500

TU(x,y)=6XY, prices are Px= 1birr and Py=2birr and income =600birr

"Budget line=Px\/PY" =1/2

"TU=6*1*2" =12

Amount of goods bought is 600/12 =50 units

The amount of each commodity is;

X=17 units

Y=33 units